# Yang-Baxter Hochschild Cohomology

### v4

@@ 5. Hopf algebra cohomology and braided algebra cohomology
where $C^{2,1}_{\rm Hf}(H,H)={\rm Hom}(H,H^{\otimes 2})$ and $C^{2,2}_{\rm Hf}(H,H)={\rm Hom}(H^{\otimes 2},H)$. If $\mu$ and $\Delta$ denote the multiplication and comultiplication of a Hopf algebra $H$, the 2-cocycles $\xi\in Z^{2,1}_{\rm Hf}(H,H)$ and $\zeta\in Z^{2,2}_{\rm Hf}(H,H)$ deform $\mu$ and $\Delta$, respectively, so that $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$ defines a Hopf algebra structure on $H\otimes{\mathbb{k}}[[\hbar]]/(\hbar^{2})$. This condition requires that $(\xi,\zeta)$ satisfies the compatibility condition needed to guarantee that $\tilde{\Delta}\tilde{\mu}=(\tilde{\mu}\otimes\tilde{\mu})(\mathbb{1}\otimes\tau\otimes\mathbb{1})(\tilde{\Delta}\otimes\tilde{\Delta})$, where $\tau$ indicates transposition of tensor factors as before.
@@
$\hskip 156.49014pt-\mu(x^{(1)}\otimes S(x^{(2)})\mu((\mathbb{1}\otimes S)\otimes\zeta(x^{(3)}))$
@@
We will denote the normalized second cohomology group of $H$ with coefficients in $H$ with the symbol $\hat{H}^{2}_{\rm Hf}(H,H):=\hat{Z}_{\rm Hf}^{2}(H,H)/(B^{1}_{\rm Hf}(H,H)\cap\hat{Z}_{\rm Hf}^{2}(H,H))$.
@@
$(\mathbb{1}\otimes\Delta)\zeta(x)=(\mathbb{1}\otimes\Delta)(x_{(1)}\otimes x_{(2)})=(x_{(1)})^{(1)}\otimes(x_{(1)})^{(2)}\otimes x_{(2)}=:\prescript{}{(1)}{x}\otimes\prescript{}{(2)}{x}\otimes\prescript{}{(3)}{x},$
and similarly we set $(\mathbb{1}\otimes\zeta)\Delta(x)=:\prescript{(1)}{}{x}\otimes\prescript{(2)}{}{x}\otimes\prescript{(3)}{}{x}$. This notation allows us to forget the brackets that determine whether the internal index is lower or upper – i.e. whether we have applied $\Delta$ or $\zeta$ first.
@@
We note that a $2$-cochain $\phi:H\otimes H\rightarrow H\otimes H$ is a YB $2$-cocycle for $R_{H}$ if and only if it is an infinitesimal deformation, i.e. if and only if $\tilde{R}=R_{H}+\hbar\phi$ is a YB operator for $H\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$. Let us consider the deformed structures $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$. They define a Hopf algebra structure (the infinitesimal deformation of $H$) since $(\xi,\zeta)$ is a Hopf $2$-cocycle by definition. Let $R_{\tilde{H}}$ be the YB operator defined through the adjoint map as in Section 2.3 for the deformation $\tilde{H}$. Since $R_{\tilde{H}}$ can be written as $R_{\tilde{H}}=R_{H}+\hbar\Psi_{\zeta}(\xi)$ we have that $\Psi_{\zeta}(\xi)$ satisfies the YB $2$-cocycle condition. Since $\xi$ is a Hochschild $2$-cocycle by definition of Hopf cohomology, to show that $(\Psi(\xi),\xi)$ is a braided algebra $2$-cocycle, we need to show that the YI and IY $2$-cocycle conditions also hold.
We apply the same argument as above to show that $(\Psi_{\zeta}(\xi),\xi)$ satisfy the YI and IY $2$-cocycle conditions. By Lemma 3.7, the deformations $\tilde{\mu}=\mu+\hbar\psi=\mu+\hbar(\xi)$ and $\tilde{R}=R+\hbar\Psi_{\zeta}(\xi)$ define a braided algebra if and only if $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle. By Lemma 2.6, if $\tilde{\mu}$ and $\tilde{\Delta}$ define a (infinitesimal) Hopf algebra, then $\mu$ and the adjoint YB operator $\tilde{R}$ define a braided algebra. Hence $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle.
We have therefore shown that $(\Psi(\xi),\xi)\in Z^{2}_{\rm YBH}(H,H)$. We now need to show that this map descends to cohomology. To do this, we need to prove that if $(\xi,\zeta)=(\delta^{1,1}_{\rm Hf}(f),\delta^{1,2}_{\rm Hf}(f))$ is cobounded and normalized, then $(\Psi_{\zeta}(\xi),\xi)$ is cobounded in YBH cohomology as well. Using the definition of Hochschild cohomology and YBH cohomology, we have that it is enough to show that $\Psi_{\zeta}(\xi)=\delta^{2}_{\rm YB}(g)$ for some $1$-cochain $g:H\rightarrow H$.
@@
$=$ $-f(x)^{(1)}\otimes f(x)^{(2)}\otimes f(x)^{(3)}+f(x^{(1)})+f(x^{(1)})\otimes x^{(2)}\otimes x^{(3)}$
$+x^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(2)}$
@@
where in the last step we have used the fact that $f(S(x))-Sf(x)+S^{\prime}(x)=0$. This shows that the map $\Psi$ induced by the correspondence $\Psi_{\zeta}(\xi)$ is well defined in cohomology.
@@ 6.2. Quadratic deformation for Hochschild cohomology
In this case we do not provide a diagrammatic proof of these classic results, as the analogue approach for braided algebras becomes increasingly complicated to handle. However, we expect that the results below that show that the obstruction for degree $2$ deformations lie in the third cohomology group of braided algebras, would extend to the case of higher degree deformations as in the case of associative algebras.
@@ 6.3. Higher order deformations for Yang-Baxter Hochschild homology
The proof consists of some straightforward, albeit long, computations. We show some details regarding the second equation. We set $R_{n+1}=\sum_{i=0}^{n+1}\hbar^{i}\phi_{i}$ and $\mu_{n+1}=\sum_{i=0}^{n+1}\hbar^{i}\mu_{i}$, as above. Since by hypothesis $(R_{n},\mu_{n})$ defines a braided algebra deformation of $(V,\mu,R)$, Equation (1) on $R_{n+1}$ and $\mu_{n+1}$ gives an obstruction only on terms of degree $n+1$ in $\hbar$. Then, writing only the terms of degree $n+1$ in $\hbar$ we have
@@
which is the second equation. Therefore, $(V,\mu_{n+1},R_{n+1})$ satisfies Equation 1 if and only if
$\delta^{2}_{\rm YI}(\phi_{n+1},\psi_{n+1})+\Xi^{\rm YI}_{n+1}-\Omega^{\rm YI}_{n+1}=0$
@@
The proof of this result is based on the use of the diagrams in Figure 12 and Figure 14, as done explicitly in [SZ-BC]. In fact, the figures can be used to compute the deviations for higher degrees to hold, from which we can infer that the obstruction to extendind a deformation to a higher degree holds up to the third differentials. Details can be found in [SZ-BC].
@@
We include the proof of Lemma 6.6 since this proof is more diagrammatic and different from methods used in [SZ-BC]. Moreover, the graphs used in the present proof were used in [SZ-BC] in a different manner.

### v5

@@ 5. Hopf algebra cohomology and braided algebra cohomology
where $C^{2,1}_{\rm Hf}(H,H)={\rm Hom}(H^{\otimes 2},H)$ and $C^{2,2}_{\rm Hf}(H,H)={\rm Hom}(H,H^{\otimes 2})$. If $\mu$ and $\Delta$ denote the multiplication and comultiplication of a Hopf algebra $H$, the 2-cocycles $\xi\in Z^{2,1}_{\rm Hf}(H,H)$ and $\zeta\in Z^{2,2}_{\rm Hf}(H,H)$ deform $\mu$ and $\Delta$, respectively, so that $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$ defines a Hopf algebra structure on $H\otimes{\mathbb{k}}[[\hbar]]/(\hbar^{2})$. This condition requires that $(\xi,\zeta)$ satisfies the compatibility condition needed to guarantee that $\tilde{\Delta}\tilde{\mu}=(\tilde{\mu}\otimes\tilde{\mu})(\mathbb{1}\otimes\tau\otimes\mathbb{1})(\tilde{\Delta}\otimes\tilde{\Delta})$, where $\tau$ indicates transposition of tensor factors as before.
@@
$\hskip 156.49014pt-\mu(x^{(1)}\otimes S(x^{(2)})\mu((\mathbb{1}\otimes S)\zeta(x^{(3)}))$
@@
We will denote the normalized second cohomology group of $H$ with coefficients in $H$ with the symbol $\hat{H}^{2}_{\rm Hf}(H,H):=\hat{Z}_{\rm Hf}^{2}(H,H)/(B^{1}_{\rm Hf}(H,H)\cap\hat{Z}_{\rm Hf}^{2}(H,H))$, where $B^{1}_{\rm Hf}(H,H)$ denotes the coboundaries in Hopf cohomology. We recall that a coboundary is a pair $(\delta^{1,1}_{\rm Hf}(f),\delta^{1,2}_{\rm Hf}(f))$ where
$\delta^{1,1}_{\rm Hf}(f)$ ::=:= $\mu(\mathbb{1}\otimes f)+\mu(f\otimes\mathbb{1})-f\mu$
$\delta^{2,1}_{\rm Hf}(f)$ ::=:= $(f\otimes\mathbb{1}+\mathbb{1}\otimes f)\Delta-\Delta f,$
for some $f:H\longrightarrow H$.
@@
$(\mathbb{1}\otimes\Delta)\zeta(x)=(\mathbb{1}\otimes\Delta)(x_{(1)}\otimes x_{(2)})=x_{(1)}\otimes(x_{(2)})^{(1)}\otimes(x_{(2)})^{(2)}=:\prescript{}{(1)}{x}\otimes\prescript{}{(2)}{x}\otimes\prescript{}{(3)}{x},$
and similarly we set $(\zeta\otimes\mathbb{1})\Delta(x)=:\prescript{(1)}{}{x}\otimes\prescript{(2)}{}{x}\otimes\prescript{(3)}{}{x}$. This notation allows us to forget the brackets that determine whether the internal index is lower or upper – i.e. whether we have applied $\Delta$ or $\zeta$ first.
@@
We note that a $2$-cochain $\phi:H\otimes H\rightarrow H\otimes H$ is a YB $2$-cocycle for $R_{H}$ if and only if it is an infinitesimal deformation, i.e. if and only if $\tilde{R}=R_{H}+\hbar\phi$ is a YB operator for $H\otimes\mathbb{k}[[\hbar]]/(\hbar^{2})$. Let us consider the deformed structures $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{\Delta}=\Delta+\hbar\zeta$. They define a Hopf algebra structure (the infinitesimal deformation of $H$) since $(\xi,\zeta)$ is a Hopf $2$-cocycle by definition. Let $R_{\tilde{H}}$ be the YB operator defined through the adjoint map as in Section 2.3 for the deformation $\tilde{H}$. Since $R_{\tilde{H}}$ can be written as $R_{\tilde{H}}=R_{H}+\hbar\Psi_{\zeta}(\xi)$ we have that $\Psi_{\zeta}(\xi)$ satisfies the YB $2$-cocycle condition. Since $\xi$ is a Hochschild $2$-cocycle by definition of Hopf cohomology, to show that $(\Psi_{\zeta}(\xi),\xi)$ is a braided algebra $2$-cocycle, we need to show that the YI and IY $2$-cocycle conditions also hold.
We apply the same argument as above to show that $(\Psi_{\zeta}(\xi),\xi)$ satisfy the YI and IY $2$-cocycle conditions. By Lemma 3.7, the deformations $\tilde{\mu}=\mu+\hbar\xi$ and $\tilde{R}=R+\hbar\Psi_{\zeta}(\xi)$ define a braided algebra if and only if $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle. By Lemma 2.6, if $\tilde{\mu}$ and $\tilde{\Delta}$ define a (infinitesimal) Hopf algebra, then $\tilde{\mu}$ and the adjoint YB operator $\tilde{R}$ define a braided algebra. Hence $(\Psi_{\zeta}(\xi),\xi)$ is a YI and IY 2-cocycle.
We have therefore shown that $(\Psi_{\zeta}(\xi),\xi)\in Z^{2}_{\rm YBH}(H,H)$. We now need to show that this map descends to cohomology. To do this, we need to prove that if $(\xi,\zeta)=(\delta^{1,1}_{\rm Hf}(f),\delta^{1,2}_{\rm Hf}(f))$ is cobounded and normalized, then $(\Psi_{\zeta}(\xi),\xi)$ is cobounded in YBH cohomology as well. Using the definition of Hochschild cohomology and YBH cohomology, we have that it is enough to show that $\Psi_{\zeta}(\xi)=\delta^{2}_{\rm YB}(f)$ for some $1$-cochain $f:H\rightarrow H$.
@@
$=$ $-f(x)^{(1)}\otimes f(x)^{(2)}\otimes f(x)^{(3)}+f(x^{(1)})\otimes x^{(2)}\otimes x^{(3)}$
$+x^{(1)}\otimes f(x^{(2)})^{(1)}\otimes f(x^{(2)})^{(2)}$
@@
where in the last step we have used the fact that $Sf(x)-f(S(x))+S^{\prime}(x)=0$. This shows that the map $\Psi$ induced by the correspondence $\Psi_{\zeta}(\xi)$ is well defined in cohomology.
@@ 6.2. Quadratic deformation for Hochschild cohomology
In this case we do not provide a diagrammatic proof of these classic results, as the analogue approach for braided algebras becomes increasingly complicated to handle. However, a different approach following the calculus of deviations of Markl-Stasheff, as described in Section 6.3, can be employed to extend the analogue of this result to the case of higher degree deformations for YBH cohomology as in the case of associative algebras. While the direct approach is not feasible at higher degrees, we still believe that it is valuable to provide a direct proof of these results of YBH cohomology at least at degree $2$. We will therefore due so, before providing the general construction.
@@ 6.3. Higher order deformations for Yang-Baxter Hochschild homology
The proof consists of some straightforward, albeit long, computations. We show some details regarding the third equation of the system. We set $R_{n+1}=\sum_{i=0}^{n+1}\hbar^{i}\phi_{i}$ and $\mu_{n+1}=\sum_{i=0}^{n+1}\hbar^{i}\mu_{i}$, as above. Since by hypothesis $(R_{n},\mu_{n})$ defines a braided algebra deformation of $(V,\mu,R)$, Equation (2) on $R_{n+1}$ and $\mu_{n+1}$ gives an obstruction only on terms of degree $n+1$ in $\hbar$. Then, writing only the terms of degree $n+1$ in $\hbar$ we have
@@
which is the third equation. Therefore, $(V,\mu_{n+1},R_{n+1})$ satisfies Equation 2 if and only if
$\delta^{2}_{\rm IY}(\phi_{n+1},\psi_{n+1})+\Xi^{\rm IY}_{n+1}-\Omega^{\rm IY}_{n+1}=0$
@@
The proof of this result is based on the use of the diagrams in Figure 12 and Figure 14, following the same approach in [SZ-BC]. In fact, the figures can be used to compute the deviations for higher degrees to hold, from which we can infer that the obstruction to extendind a deformation to a higher degree holds up to the third differentials. Details can be found in [SZ-BC].
@@
We include the proof of Lemma 6.6 in Appendix B below since this proof is more diagrammatic and different from methods used in [SZ-BC]. Moreover, the graphs used in the present proof were used in [SZ-BC] in a different manner.